A normal distribution has a mean equal to 48. What is the standard deviation of this normal distribution if 2.5% of the proportion under the curve lies to the right of x = 55.84? (Round your answer to two decimal places.)
Answer #1

Given that,

P(Z > 1.96) = 0.025

z = 1.96

Using z-score formula,

= (x - ) / z = (55.84 - 48) / 1.96 = 4

The standard deviation of this normal distribution = 4